OpenStudy (anonymous):

3/a^-9b A.27a/b B. 3/a^9b C. 3/ab^-8 D. 3a^9/b

OpenStudy (hoblos):

3/a^-9b = (3/b)(1/a^-9) = (3/b)(a^9) = 3a^9/b